26
In an urn I have 20 marbles: 2 red, 3 yellow, 4 blue, 5 green, and 6 black marbles. The selection of one marble is random. The winning table is below. What is the expected value of winnings?
Color Red Yellow Blue Green Black
Amount won $10 $5 $2 $1 $0

Answer by greenestamps(13334)



Expected value is the sum the products for each color of the probability of getting that color and the payout for that color.

(2/20)*$10 + (3/20)*$5 + (4/20)*$2 + (5/20)*$1 + (6/20)*$0 = (20+15+8+5+0)/20 = 48/20 = $2.40.